OpenStudy (anonymous):

Find the half-life for a substance that decays by 7% each year.

OpenStudy (anonymous):

y=a(1+(-7%))^t let retriceume initial amount R.A.I.X:100g half = 50 y=a(1+(-0.07))^t y=a(1+(0.93)^t (50/100)=(100*(0.93))^t/(100) (1/2)=(0.93)^t ( ln(1/2)/(ln(0.93)=t (-0.693147)/(-0.072570)=t t=9.56 years